# Integrating Graceful Degradation and Recovery through Requirement-driven Adaptation

### 18 Jan 2024 v1

@@ IX Limitation and Future Work
We propose a requirement-driven runtime adaptation framework that coordinates between grace degradation and recovery. Through the case studies on UUVs, we have demonstrated our proposed approach can result in more desirable system behaviors and higher satisfaction with system requirements.
@@
However, our approach does have several assumptions about the system that it deploys to which may limit its applicability. First, we assume the system has a meaningful way of quantifying the satisfaction of its requirements, and that this requirement can be weakenable, and it is acceptable for the user to have temporary degradation to the system performance. Thus, this framework cannot be applied to safety-critical requirements where the requirements are hard constraints and cannot be negotiated, (i.e. collision avoidance system for smart vehicles, or Airborne Collision Avoidance System (ACAS) for commercial aircraft).
@@
Secondly, we assume that the behavior of the system and the interaction between the system and the external environment (i.e. how the vehicles move around the water) can be predicted by a deterministic environmental model, which is used for requirement analysis and planning. However, the estimation from the environmental model may deviate from reality, due to environmental disturbance and other uncertainties. To address this, we used model predictive control [6], which involves repeated prediction and planning on a short horizon to reduce the effect of deviation in the estimation.
Thirdly, we cannot guarantee that a MILP solver can return a satisfiable solution. This can be due to over-restrictive environmental models, or a limited range of alternative requirements that the solver can search for. In our case studies, we cope with this problem by implementing a fallback adaptation mechanism (i.e. revert to pre-defined adaptation actions instead of dynamically generated actions), and extensive tuning of the environmental model and range of alternative requirements.
Finally, besides using requirement-driven adaptation to coordinate between graceful degradation and recovery, other methods can be used to coordinate them in more interesting ways. For example, we can use probabilistic-based analysis to determine how the system can be degraded by predicting the cost and feasibility of recovery.

### 08 Apr 2024 v2

@@ 9. Limitations and Future Work
We propose a requirement-driven runtime adaptation framework that coordinates between grace degradation and recovery. Through the case study on UUVs, we have demonstrated our proposed approach can result in more desirable system behaviors during periods of degradation and recovery.
However, our approach makes several assumptions about the system that may limit its applicability. First, our approach tackles the class of requirements can be specified using STL (Maler and Nickovic, 2004). These requirements are specific to cyber-physical systems, where the system is time-sensitive, and sensor inputs can be monitored and transformed into continuous signals and quantified using state predicates. Thus, the framework is currently not designed to handle other classes of requirements (e.g., discrete behavioral specifications in LTL or stochastic ones in PCTL(Ciesinski and Größer, 2004)).
@@
Secondly, we assume the system has a meaningful way of quantifying the satisfaction of its requirements, these requirements can be weakened, and the user is willing to tolerate temporary degradation in the associated system utility. Thus, this framework is not well-suited for safety-critical requirements, which tend to be hard constraints and cannot be negotiated (i.e., a collision avoidance system for smart vehicles).
@@
Thirdly, we assume that interactions between the system and the external environment (i.e. how a UUV move around the water based on an actuator command) can be predicted by a deterministic environmental model. However, the estimation from the environmental model may deviate from reality, due to environmental disturbance and other uncertainties. To address this, we used model predictive control (Rawlings and Mayne, 2009), which involves repeated prediction and planning on a short horizon to reduce the effect of deviation in the estimation.
Another limitation of our approach is that a MILP solver may fail to find a satisfiable solution. This can be due to an overly restrictive environmental model, a limited range of alternative requirements that the solver can search for, or an unresolvable system state at the time of the adaptation. While this happened rarely in our case studies (less than 2% of adaptation cycles), we can cope with the problem through some minor tuning of the adaptation frameworks like reducing constraints to the environmental model, increasing the ranges of requirements available, or defining some simple fallback adaptation approaches, etc.
In future work, we also plan to investigate the applicability of our approach in our domains, such as robotics and cyber-security, where the ability to gracefully degrade and recover is critical for system autonomy and resilience.
10. Acknowledgement
This work was supported in part by Award N00014172899 from the Office of Naval Research, Award No. H98230-23-C-0274 from the National Security Agency, and the National Science Foundation award CCF-2144860. Any views, opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agencies.